Find the distance and angle of a vector if the x-component is 38.37 m and the y-component is 61.30m. 


1
Expert's answer
2022-02-05T10:28:22-0500
"|r|=\sqrt{x^2+y^2}=72.32\text{ m}."

"\theta=\arctan(y/x)=57.96°."